# The stability of propagating plane inertial waves in rotating fluids

### v1

@@ III.1 Verification of Linear Theory with DNS in the Comoving Frame
The anisotropy of the wavenumber orientation of the most unstable perturbations can also be confirmed in real-space visualizations of the flow. Figure 6 shows the vorticity of the perturbations along the rotation vector $\bm{\Omega}^{\prime}\cdot\bm{\omega}^{\prime}$ (where $\bm{\omega}^{\prime}=\nabla\times\bm{u}^{\prime}$) for the three fiducial simulations. One can see that the perturbations disturbing the base flow of the low frequency wave $\omega/2\Omega=0.17$ have small structures along the $\hat{\bm{y}}^{\prime}$ direction, with extended structures that are nearly invariant in the $\hat{\bm{x}}^{\prime}$ direction. Perturbations of the intermediate frequency wave $\omega/2\Omega=0.71$ also have variations primarily in the $\hat{\bm{y}}^{\prime}$ direction, but also have significant structure in the $\hat{\bm{x}}^{\prime}$ direction. By contrast, there is no obvious anisotropy in the perturbations of the high frequency wave with $\omega/2\Omega=0.98$.
@@ III.2 Nonlinear Evolution in the Rotating Frame
As an IW propagates, unstable perturbations grow at the expense of the wave’s energy, eventually disrupting the flow of the IW and draining its energy. In a rotating fluid, the energy can be converted into two qualitatively different channels. One channel is a forward cascade on dynamical timescales $\sim 1/A^{\prime}|\omega|$ wherein nonlinear interactions progressively transfer energy to smaller scales, ultimately converting the energy into heat through viscous dissipation. The other channel is conversion of the energy into geostrophic modes, which are 2D flows invariant along the rotation axis, with $\bm{\Omega}\cdot\nabla\bm{u}=0$. This conversion in our model must occur through near-resonant triadic or resonant quartic interactions, since exact triadic resonances cannot transfer energy from IWs to geostrophic modes due to Greenspan’s theorem [e.g. 47, 1]. The geostrophic motions then inverse cascade through mergers of 2D vortices and decay on long viscous timescales due to their lack of variation in the vertical direction and large horizontal scales [e.g. 4]. Below, we explore the evolution of the flow and the partition of the initial wave energy between the geostrophic flow and the forward cascade after the primary IW has broken down.
@@
Simulations of propagating IWs are carried through the nonlinear phase of their evolution in the rotating frame. A visualization of a representative simulation is shown in Figure 7 for a primary IW with $A^{\prime}=0.3$ and $\omega/2\Omega=0.17$. The flow evolution is split into three distinct phases. First, the linear phase occurs at early times $\omega t\lesssim 180$ when the unstable perturbations grow exponentially (not shown). The viscous damping of the primary wave during this phase is seen as the slow decrease of the IW energy $E_{\rm IW}=0.5|\bm{u}_{\bm{k},\rm IW}|^{2}$ in panel (a). At the end of the linear phase, the unstable perturbations develop significant amplitudes, as seen in panel (b). They become nonlinear and saturate during a turbulent phase $180\lesssim\omega t\lesssim 500$ when the primary wave energy is drained. The primary wave flow is disrupted early on during this phase, as seen in panel (c) at $\omega t=210$ when the flow has both small scale features characteristic of a forward cascade and vertically-stretched features characteristic of geostrophic motions. The energy in the geostrophic modes $E_{\rm gs}=\sum_{\bm{k}\;(k_{z}=0)}|u_{\bm{k}}|^{2}$ grows rapidly until $\omega t\sim 220$ when the primary wave energy is fully drained. Afterwards, $E_{\rm gs}$ remains nearly constant while the energy in the non-geostrophic flow $E_{\rm tot}-E_{\rm gs}$ (the separation between the blue and green curves) decays through a forward cascade on longer timescales until $\omega t\sim 500$. The last phase of evolution occurs for $\omega t\gtrsim 500$ when $E_{\rm gs}\approx E_{\rm tot}$ is dominated by the dynamics of merging geostrophic vortices, as is seen in panel (c) at $\omega t=600$. Eventually, the vortices will reach the box scale and decay on slow, viscous timescales. Similar behavior has been observed in simulations of the elliptical and precessional instabilities [e.g. 48, 49, 50, 51], which continually excite IW unlike the decaying case here.
The three phases of evolution can also be examined in the 1D horizontal energy spectra $E(k_{\rm h})$ shown in Figure 8, where $k_{\rm h}=\sqrt{k_{x}^{2}+k_{y}^{2}}$. At early times $\omega t\lesssim 180$, the spectrum has peak at $k_{\rm h}L/2\pi\sim 6$ corresponding to the energy in the primary wave. The energy at other wavenumbers is determined by the growth rates of the unstable perturbations. As the unstable perturbations saturate and the wave breaks down, the spectrum begins to flatten out during the most turbulent part of the flow evolution, which extracts almost all of the primary wave energy within a few wave periods. However, the spectrum remains far from a Kolmogorov scaling $E(k_{\rm h})\propto k_{\rm h}^{-5/3}$. At later times, the energy spectrum approaches the $E(k_{\rm h})\propto k_{\rm h}^{-3}$ scaling expected for geostrophic turbulence [e.g. 47], which appears as a flat compensated spectrum.
To examine how the nonlinear saturation depends on the properties of the primary IW, we run a set of simulations varying the frequency in a range $0.17\leq\omega/2\Omega\leq 0.98$ for two amplitudes $A^{\prime}=0.3,1.0$. The resulting fraction $E_{\rm gs}/E_{0}$ of the initial wave energy $E_{0}$ trapped in the geostrophic component of the flow after wave breakdown is compiled in Figure 9. The measurement of $E_{\rm gs}/E_{0}$ is done at the end of every simulation when $E_{\rm gs}$ contains at least $95\%$ of the total energy remaining in the flow. One can see that lower frequency waves ($|\omega|\lesssim\Omega$) tend to have a larger portion of their energy transferred into the geostrophic flow, upwards of $E_{\rm gs}/E_{0}\gtrsim 10\%$. On the other hand, higher frequency waves ($|\omega|\gtrsim\Omega$) lose most of their energy to the forward cascade, with only $E_{\rm gs}/E_{0}\lesssim 5\%$ in the geostrophic flow. This trend may be due to the higher efficiency of nearly-horizontal near-resonance triads $(\omega_{p}+\omega_{q}+\omega_{r}={\cal O}(A^{\prime}))$ with $\omega_{p}/2\Omega={\cal O}(A^{\prime})$, $\omega_{q}/2\Omega={\cal O}(A^{\prime})$, $\omega_{r}=0$, in transferring energy into the geostrophic mode [47], where the low frequency primary wave participates as one of the nearly-geostrophic modes. We caution that part of the non-monotonic trend of $E_{\rm gs}/E_{0}$ with $\omega$ for $A^{\prime}=1.0$ at lower frequencies (where the wavenumber of the primary wave is highly inclined) may be due to box-size effects, as the number of (near) resonant triads available for interactions may be limited by the discreteness of the numerical Fourier grid.
The conversion fraction $E_{\rm gs}/E_{0}$ also depends on the wave amplitude. Lower amplitude waves tend to have a smaller conversion fraction, at least for higher frequency waves with $\omega/2\Omega\gtrsim 0.3$. This trend is consistent with some of the proposed mechanisms that can transfer energy into geostrophic modes with a rate $\sigma_{\rm gs}\sim A^{\prime 2}\omega$ when $A^{\prime}\lesssim 1$, which include near-resonant triadic [52, 47, 53, 54, 55] and resonant four-mode interactions [52, 35, 47, 56]. These mechanisms have been used to explain the steady-state fraction of energy in the geostrophic component of forced rotating turbulence, and we expect that these same mechanisms operate during the breakdown of a single IW. The scaling $\sigma_{\rm gs}/\sigma_{\rm PSI}\propto A^{\prime}$ suggests that at sufficiently low wave amplitudes, the pumping of geostrophic modes will be too slow to meaningfully operate within the timescale for wave breakdown $\sim 1/\sigma_{\rm PSI}$. However, we leave exploration of the saturation of lower amplitude waves with $A^{\prime}\lesssim 0.1$ for future work due to the increased computational cost of simulating flows with low Ekman numbers, including the longer run time required for the instabilities to grow and saturate.
@@
Figure 7: Simulation in the rotating frame of a propagating IW with amplitude $A^{\prime}=0.3$ and frequency $\omega/2\Omega=0.17$ (angle $\theta=80^{\circ}$). Panel (a) shows the temporal evolution of the total energy $E_{\rm tot}$ of the flow field and the contributions from the energy in the primary wave $E_{\rm IW}$ and in the geostrophic flow $E_{\rm gs}$. Panels (b-d) show the vertical vorticity $\omega_{z}=(\nabla\times\bm{u})\cdot\hat{\bm{z}}$ (the component aligned with $\bm{\Omega}=\Omega\hat{\bm{z}}$) at three times representing the linear, turbulent, and decay phases of the flow evolution.
@@
Figure 8: Energy spectra of the velocity field at representative times (colored coded) during the evolution of a propagating IW. The simulation is the same as the one shown in Figure 7. Panel (a) shows the horizontal energy spectra $E(k_{\rm h})$ compensated by $k_{\rm h}^{3}$ and panel (b) shows the vertical energy spectra $E(k_{z})$ compensated by $k_{z}^{3}$. The compensated horizontal energy spectrum is nearly flat at late times when the flow is nearly geostrophic.
@@ IV Conclusion
Our analysis combined a comprehensive numerical Floquet method (complemented with an analytical asymptotic analysis) with direct numerical simulations to investigate the linear stability and nonlinear saturation of propagating plane inertial waves. We found that the fastest-growing unstable perturbations have a non-dimensional growth rate $\sigma^{\max}/(A\,k)\approx 0.3$ that is nearly independent of the wave amplitude, where $A$ is the velocity amplitude and $k$ is the wavenumber. This rate closely matches the growth rate of the parametric subharmonic instability at low amplitudes $A\ll|\omega|/k$ (where $\sigma_{\mathrm{PSI}}/(A\,k)\approx 0.325$) and decreases only slightly by around $\sim 15\%$ at larger amplitudes $A\sim|\omega|/k$. We also found that the most unstable wavenumbers are much larger than the wavenumber of the primary IW at low wave amplitudes, as expected for the PSI, but become comparable with the primary’s wavenumber at high amplitudes. Furthermore, the orientation of the most unstable wavenumbers strongly depends on the primary IW frequency. For low-frequency waves $|\omega|\ll 2\Omega$, the perturbations are aligned orthogonal to the plane of the rotation axis and the primary IW wavevector, while for high-frequency waves $|\omega|\lesssim 2\Omega$, their distribution becomes nearly isotropic. This suggests that a single IW is generally disrupted by perturbations that isotropize the flow in the plane perpendicular to $\bm{\Omega}$ as the wave breaks down, in close agreement with the findings of [39] and [40].
@@
We then used DNS to study the breakdown of propagating IWs in a triply-periodic domain. Our simulations first cross-validated the Floquet theory and then explored the nonlinear evolution of the flow. When the primary IW breaks down, its energy can be transferred into two complementary channels. It can either be dissipated through a forward cascade on dynamical timescales or accumulated in slowly evolving geostrophic modes. The relative efficiency of these two channels depends on the interactions between the primary IW and the saturating unstable perturbations. We found that lower-frequency waves generally convert a larger fraction of initial wave energy into geostrophic motions (e.g., a fraction of more than $10\%$ for IWs with $|\omega|\lesssim\Omega$ and high amplitudes with $A\sim|\omega|/k$). In contrast, higher-frequency waves lose most of their energy through the forward cascade. The conversion fraction to geostrophic energy also decreases as the wave amplitude decreases, with the exception of the lowest-frequency waves.
@@
The turbulence resulting from breakdown of a single IW may share aspects of forced rotating turbulence, which pumps geostrophic modes when the Rossby number is above a threshold value. As geostrophic modes cannot gain energy from triadic resonance interactions in this triply-periodic model, the accumulation of energy in geostrophic modes likely occurs through near-resonant triadic interactions [52, 47, 53, 54, 55] or higher-order resonant interactions [52, 35, 47, 56], which become less efficient at lower Rossby numbers. We suggest that studying the breakdown of IWs could be a cleaner numerical experiment for understanding these processes. Discerning the threshold Rossby number (wave amplitude) and the mechanisms for pumping geostrophic modes during breakdown of a single IW requires future work.
@@
The linear stability problem and simulations can also be extended in future work to include more realistic backgrounds, such as stable stratification, differential rotation, or magnetic fields. Stratification is particularly important in oceanic environments where IWs are excited predominately by wind stresses and are a major source of mixing in the upper oceans [5], impacting oceanic biogeochemistry and the climate [18]. In astrophysical contexts, IWs excited by planet-star tidal interactions can propagate in the convection zones of their host stars (or inside giant planets themselves) where differential rotation and/or large-scale magnetic fields modify wave propagation and instability [e.g. 13, 8, 9, 57]. In close orbits or when excited in resonance, IWs are expected to achieve large amplitudes, and understanding their general propagation, stability, and dissipation is key for predicting tidal evolution.
@@
We note that in realistic physical systems, either in experimental or natural fluids, IWs are typically excited in beams of finite width. The linear instability analysis in our triply-periodic model is applicable when unstable daughter waves, whose group velocities would be misaligned with the primary wave beam, have a growth timescale that is shorter than the time it takes them to travel out of the beam. It is well known that finite-size effects in IGW beams restrict the wavenumber space of unstable modes that can cause wave breakdown [58, 59, 60, 61, 62] and thus affects how far an IGW beam propagates before disrupting and depositing its energy and momentum. Our results for the linear stability of IWs can be similarly extended to understand the propagation of finite-amplitude IW beams and their transport properties.

### v2

@@ III.1 Verification of Linear Theory with DNS in the Comoving Frame
The anisotropy of the wave vector orientation of the most unstable perturbations can also be confirmed in real-space visualizations of the flow. Figure 7 shows the vorticity of the perturbations along the rotation vector $\bm{\Omega}^{\prime}\cdot\bm{\omega}^{\prime}$ (where $\bm{\omega}^{\prime}=\nabla\times\bm{u}^{\prime}$) for the three fiducial simulations. One can see that the perturbations disturbing the base flow of the low frequency wave $\omega/2\Omega=0.17$ have small structures along the $\hat{\bm{y}}^{\prime}$ direction, with extended structures that are nearly invariant in the $\hat{\bm{x}}^{\prime}$ direction. Perturbations of the intermediate frequency wave $\omega/2\Omega=0.71$ also have variations primarily in the $\hat{\bm{y}}^{\prime}$ direction, but also have significant structure in the $\hat{\bm{x}}^{\prime}$ direction. By contrast, there is no obvious anisotropy in the perturbations of the high frequency wave with $\omega/2\Omega=0.98$.
@@ III.2 Nonlinear Evolution in the Rotating Frame
As an IW propagates, unstable perturbations grow at the expense of the wave’s energy, eventually disrupting the flow of the IW and draining its energy. In a rotating fluid, the energy can be converted into two qualitatively different channels. One channel is a forward cascade on dynamical timescales $\sim 1/A^{\prime}|\omega|$ wherein nonlinear interactions progressively transfer energy to smaller scales, ultimately converting the energy into heat through viscous dissipation. The other channel is conversion of the energy into geostrophic modes, which are 2D flows invariant along the rotation axis, with $\bm{\Omega}\cdot\nabla\bm{u}=0$. This conversion in the triply periodic model must occur through near-resonant triadic or resonant quartic interactions, since exact triadic resonances cannot transfer energy from IWs to geostrophic modes due to Greenspan’s theorem [e.g. 60, 34]. The geostrophic motions then inverse cascade through mergers of 2D vortices and decay on long viscous timescales due to their lack of variation in the vertical direction and large horizontal scales [e.g. 32]. Below, we explore the evolution of the flow and the partition of the initial wave energy between the geostrophic flow and the forward cascade after the primary IW has broken down.
@@
Simulations of propagating IWs are carried out through the nonlinear phase of their evolution in the rotating frame. A visualization of a representative simulation is shown in Figure 8 for a primary IW with $A^{\prime}=0.3$ and $\omega/2\Omega=0.17$. The flow evolution is split into three distinct phases. First, the linear phase occurs at early times $\omega t\lesssim 180$ when the unstable perturbations grow exponentially (not shown). The viscous damping of the primary wave during this early period is seen as the slow, exponential decay of the IW energy $E_{\rm IW}=0.5|\bm{u}_{\bm{k},\rm IW}|^{2}$ in panel (a). As the unstable perturbations develop significant amplitudes (panel (b) at $\omega t=180$), the flow enters the second, highly turbulent phase $180\lesssim\omega t\lesssim 220$ when the unstable modes saturate. This phase is characterized by drainage of the primary wave energy, accompanied by rapid growth of energy in the geostrophic modes $E_{\rm gs}=\sum_{\bm{k}\;(k_{z}=0)}|u_{\bm{k}}|^{2}$. The flow during this phase has both small-scale features characteristic of a forward cascade and vertically stretched features characteristic of geostrophic motions (panel (c) at $\omega t=210$). The flow then transitions into a third, decay phase where $E_{\rm gs}$ remains nearly constant while the energy in the non-geostrophic flow $E_{\rm tot}-E_{\rm gs}$ (the separation between the blue and green curves in panel (a)) decays through a forward cascade on longer timescales. This phase becomes dominated by the dynamics of merging geostrophic vortices once $E_{\rm gs}\approx E_{\rm tot}$ (panel (d) at $\omega t=600$). Eventually, the vortices will reach the box scale and decay on slow, viscous timescales. Similar behavior is observed in simulations of the elliptical and precessional instabilities [e.g. 10, 12, 57, 26], which continually excite IWs, unlike the decaying case here.
@@
Figure 8: Simulation in the rotating frame of a propagating IW with amplitude $A^{\prime}=0.3$, frequency $\omega/2\Omega=0.17$ (angle $\theta=80^{\circ}$), and Ekman number $\mathrm{Ek}=10^{-3}$. Panel (a) shows the temporal evolution of the total energy $E_{\rm tot}$ of the flow field and the contributions from the energy in the primary wave $E_{\rm IW}$ and in the geostrophic flow $E_{\rm gs}$. The dashed line shows $E_{\rm IW}(t)$ if the primary wave decays only by viscous damping $(\propto\exp(-2\mathrm{Ek}k_{\mathrm{IW}}^{2}t))$. Panels (b-d) show the vertical vorticity $\omega_{z}=(\nabla\times\bm{u})\cdot\hat{\bm{z}}$ (the component aligned with $\bm{\Omega}=\Omega\hat{\bm{z}}$) at three times representing the linear, turbulent, and decay phases of the flow evolution.
@@
Figure 9: Energy spectra of the velocity field at representative times (color-coded) during the evolution of a propagating IW. The simulation is the same as the one shown in Figure 8. Panel (a) shows the horizontal energy spectra $E(k_{\rm h})$ compensated by $k_{\rm h}^{3}$ and panel (b) shows the vertical energy spectra $E(k_{z})$ compensated by $k_{z}^{3}$. The compensated horizontal energy spectrum at intermediate scales is nearly flat at late times when the flow is nearly geostrophic.
The three evolutionary phases can also be examined in the 1D horizontal and vertical energy spectra, $E(k_{\rm h})$ and $E(k_{z})$, shown in Figure 9, where $k_{\rm h}=\sqrt{k_{x}^{2}+k_{y}^{2}}$. The initial horizontal and vertical spectra at $t=0$ have peaks at $k_{\rm h}L/2\pi\sim 6$ and $k_{z}L/2\pi\sim 1$, respectively, corresponding to the primary wave. They also contain power law backgrounds $E(k_{\mathrm{h}})\propto k_{\mathrm{h}}$ and $E(k_{z})\propto k_{z}^{0}$ corresponding to the solenoidal Gaussian noise velocity field. At early times $\omega t\lesssim 180$, the energy at the peaks remains unchanged while the energy at other wave vectors is determined by the growth rates of the unstable perturbations (the largest wavenumbers are viscously damped). As the unstable perturbations saturate and the wave breaks down, the horizontal spectrum begins to flatten out during the most turbulent part of the flow’s evolution. However, it remains far from a Kolmogorov scaling $E(k_{\rm h})\propto k_{\rm h}^{-5/3}$. At later times, the horizontal energy spectrum approaches the $E(k_{\rm h})\propto k_{\rm h}^{-3}$ scaling expected for geostrophic turbulence [e.g. 60], which appears as a flat compensated spectrum at intermediate scales. The geostrophic-nature of the flow is evidenced in the vertical spectrum by the accumulation of energy in the $k_{z}=0$ modes at late times (the bin at $k_{z}L/2\pi=0.5$ in panel (b) holds the energy of modes with $k_{z}L/2\pi\leq 0.5$).
To examine how the nonlinear saturation depends on the properties of the primary IW, we run a set of simulations varying the frequency in a range $0.17\leq\omega/2\Omega\leq 0.98$ for two amplitudes $A^{\prime}=0.3,1.0$. The resulting fraction $E_{\rm gs}/E_{0}$ of the initial wave energy $E_{0}$ trapped in the geostrophic component of the flow after wave breakdown is compiled in Figure 10. The measurement of $E_{\rm gs}/E_{0}$ is done at the end of every simulation when $E_{\rm gs}$ contains at least $95\%$ of the total energy remaining in the flow. One can see that lower frequency waves ($|\omega|\lesssim\Omega$) tend to have a larger portion of their energy transferred into the geostrophic flow, upwards of $E_{\rm gs}/E_{0}\gtrsim 10\%$. On the other hand, higher frequency waves ($|\omega|\gtrsim\Omega$) lose most of their energy to the forward cascade, with only $E_{\rm gs}/E_{0}\lesssim 5\%$ in the geostrophic flow. This trend may be due to the higher efficiency of nearly-horizontal near-resonance triads $(\omega_{p}+\omega_{q}+\omega_{r}={\cal O}(A^{\prime}))$ with $\omega_{p}/2\Omega={\cal O}(A^{\prime})$, $\omega_{q}/2\Omega={\cal O}(A^{\prime})$, $\omega_{r}=0$, in transferring energy into the geostrophic mode [60], where the low frequency primary wave participates as one of the nearly-geostrophic modes. We caution that part of the non-monotonic trend of $E_{\rm gs}/E_{0}$ with $\omega$ for $A^{\prime}=1.0$ at lower frequencies (where the wave vector of the primary wave is highly inclined) may be due to box-size effects, as the number of (near) resonant triads available for interactions may be limited by the discreteness of the numerical Fourier grid.
The conversion fraction $E_{\rm gs}/E_{0}$ also depends on the wave amplitude. Lower amplitude waves tend to have a smaller conversion fraction, at least for higher frequency waves with $\omega/2\Omega\gtrsim 0.3$. This trend is consistent with some of the proposed mechanisms that can transfer energy into geostrophic modes with a rate $\sigma_{\rm gs}\sim A^{\prime 2}|\omega|$ when $A^{\prime}\lesssim 1$, which include near-resonant triadic [53, 60, 59, 27, 45] and resonant four-mode interactions [53, 40, 60, 21]. These mechanisms have been used to explain the steady-state fraction of energy in the geostrophic component of forced rotating turbulence, and we expect that these same mechanisms operate during the breakdown of a single IW. The scaling $\sigma_{\rm gs}/\sigma_{\rm PSI}\propto A^{\prime}$ suggests that at sufficiently low wave amplitudes, the pumping of geostrophic modes will be too slow to meaningfully operate within the timescale for wave breakdown $\sim 1/\sigma_{\rm PSI}$. However, we leave exploration of the saturation of lower amplitude waves with $A^{\prime}\lesssim 0.1$ for future work due to the increased computational cost of simulating flows with low Ekman numbers, including the longer run time required for the instabilities to grow and saturate.
Lastly, we note that the conversion fraction $E_{\rm gs}/E_{0}$ may also depend on the choice of initial perturbations, including their amplitude and distribution in wave vector space. Larger amplitude noise reduces the amount that the primary wave decays due to viscosity because unstable modes saturate earlier. It also allows slower growing modes to participate in the saturation phase, in contrast to the case of extremely small noise amplitude where mainly the fastest growing modes contribute. Despite these effects, we have found that the final value of $E_{\rm gs}/E_{0}$ is weakly altered for large noise amplitudes, changing by at most a few percent for initial noise energy up to $(E_{\rm tot}-E_{0})/E_{\rm tot}\sim 0.01$. We leave exploring different noise spectra, or possibly seeding single unstable eigenmodes, for future work.
@@ IV Conclusion
Our analysis combined a comprehensive numerical Floquet method (complemented with an analytical asymptotic analysis) with direct numerical simulations to investigate the linear stability and nonlinear saturation of propagating plane IWs. We found that the fastest-growing unstable perturbations have a non-dimensional growth rate $\sigma^{\max}/(A\,k)\approx 0.3$ that is nearly independent of the wave amplitude, where $A$ is the velocity amplitude and $k$ is the wavenumber. This rate closely matches the growth rate of the PSI at low amplitudes $A\ll|\omega|/k$ (where $\sigma_{\mathrm{PSI}}/(A\,k)\approx 0.325$) and decreases only slightly by around $\sim 15\%$ at larger amplitudes $A\sim|\omega|/k$. We also found that the most unstable wavenumbers are much larger than the wavenumber of the primary IW at low wave amplitudes, as expected for the PSI, but become comparable with the primary’s wavenumber at high amplitudes. Furthermore, the orientation of the wave vector of the most unstable modes strongly depends on the primary IW frequency. For low-frequency waves $|\omega|\ll 2\Omega$, the wave vectors are aligned orthogonal to the plane of the rotation axis and the primary IW wavevector, while for high-frequency waves $|\omega|\lesssim 2\Omega$, their distribution becomes nearly isotropic. This suggests that a single IW is generally disrupted by perturbations that isotropize the flow in the plane perpendicular to $\bm{\Omega}$ as the wave breaks down, in close agreement with the findings of [52] and [1].
@@
We then used DNS to study the breakdown of propagating IWs in a triply periodic domain. Our simulations first cross-validated the Floquet theory and then explored the nonlinear evolution of the flow. When the primary IW breaks down, its energy can be transferred into two complementary channels. It can either be dissipated through a forward cascade on dynamical timescales or accumulated in slowly evolving geostrophic modes. The relative efficiency of these two channels depends on the interactions between the primary IW and the saturating unstable perturbations. We found that lower-frequency waves generally convert a larger fraction of initial wave energy into geostrophic motions (e.g., a fraction of more than $10\%$ for IWs with $|\omega|\lesssim\Omega$ and high amplitudes with $A\sim|\omega|/k$). In contrast, higher-frequency waves lose most of their energy through the forward cascade. The conversion fraction to geostrophic energy also decreases as the wave amplitude decreases, with the exception of the lowest-frequency waves.
@@
The turbulence resulting from the breakdown of a single IW may share aspects of forced rotating turbulence, which pumps geostrophic modes when the Rossby number is above a threshold value. As geostrophic modes cannot gain energy from triadic resonance interactions in this triply periodic model, energy accumulation in geostrophic modes likely occurs through near-resonant triadic interactions [53, 60, 59, 27, 45] or higher-order resonant interactions [53, 40, 60, 21], which are less efficient at lower Rossby numbers. We suggest that studying the breakdown of IWs could be a cleaner numerical experiment for understanding these processes. Discerning the threshold Rossby number (wave amplitude) and the mechanisms for pumping geostrophic modes during the breakdown of a single IW requires future work.
@@
The linear stability problem and simulations can also be extended to include more realistic backgrounds, such as stable stratification, differential rotation, or magnetic fields. Stratification is particularly important in oceanic environments where IWs are excited predominantly by wind stresses and are a major source of mixing in the upper oceans [3], impacting oceanic biogeochemistry and the climate [37]. In astrophysical contexts, IWs excited by planet-star tidal interactions can propagate in the convection zones of their host stars (or inside giant planets themselves) where differential rotation and/or large-scale magnetic fields modify wave propagation and instability [e.g. 55, 5, 6, 7]. In close orbits or when excited in resonance, IWs are expected to achieve large amplitudes, and understanding their general propagation, stability, and dissipation is key for predicting tidal evolution.
@@
We note that in realistic physical systems, either in experimental or natural fluids, IWs are typically excited in beams of finite width. The linear instability analysis in our triply periodic model is applicable when unstable daughter waves, whose group velocities would be misaligned with the primary wave beam, have a growth timescale that is shorter than the time it takes them to travel out of the beam. It is well known that finite-size effects in IGW beams restrict the wave vector space of unstable modes that can cause wave breakdown [43, 18, 19, 24, 33] and thus affect how far an IGW beam propagates before disrupting and depositing its energy and momentum. Our results for the linear stability of IWs can be similarly extended to understand the propagation and transport properties of finite-amplitude IW beams.